# Localized RETE for Incremental Graph Queries with Nested Graph Conditions

### v1

@@ A.5. Supplementary Lemmata
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ consists of $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$, $(N^{sat}_{(Q^{\prime},\psi^{\prime})},p^{sat}_{(Q^{\prime},\psi^{\prime})})=localize^{sat}(Q^{\prime},\psi^{\prime})$, $[\cup]^{\Phi}$, $RPS_{r}=RPS([\cup]^{\Phi},N^{\Phi}_{Q})$, and $p^{sat}=[\ltimes]^{\Phi}$. The execution order for $(N^{sat},p^{sat})$ is then given by $order^{sat}(N^{sat})=order(N^{\Phi}_{Q^{\prime}})\circ order^{sat}(N^{sat}_{(Q^{\prime},\psi^{\prime})})\circ[\cup]^{\Phi}\circ toposort(RPS_{r})^{-1}\circ order(N^{\Phi}_{Q})\circ[\ltimes]^{\Phi}$.
By Lemma 13 in [BG24b], $\mathcal{C}_{1}^{\Phi}=execute(order(N^{\Phi}_{Q^{\prime}}),N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}_{Q^{\prime}}$. $\mathcal{C}_{2}^{\Phi}=execute(order^{sat}(N^{sat}_{(Q^{\prime},\psi^{\prime})}),N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{1})$ is then also consistent for all nodes in $N^{sat}_{Q^{\prime}}$ by the induction hypothesis. By Lemma 13 in [BG24b], $\mathcal{C}^{\Phi}=execute([\cup]^{\Phi}\circ toposort(RPS_{r})^{-1}\circ order(N^{\Phi}_{Q})\circ[\ltimes]^{\Phi},N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{2})$ must then be consistent for all nodes in $N^{sat}$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. It then holds that $\sum_{n^{sat}\in V^{N^{sat}}}|\mathcal{C}^{\Phi}(n^{sat})|\leq 25\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@
For each nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$, $N^{sat}$ contains the subnets $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$ and $(N^{\Phi}_{Q^{\prime\prime}},p^{\Phi}_{Q^{\prime\prime}})=localize(Q^{\prime\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$, a marking-sensitive semi-join $[\ltimes]^{\Phi}$, as well as the request projection structure $RPS_{l}=RPS([\cup]^{\Phi},N^{\Phi}_{Q^{\prime}})$. $N$ contains a RETE subnet $(N_{Q^{\prime}},p_{Q^{\prime}})$ computing matches for $Q^{\prime}$ and a semi-join []delimited-[]left-normal-factor-semidirect-product[$\ltimes$][ ⋉ ].
By Lemma 38 and because the additional dependency on $RPS_{l}$ does not affect the worst case of a fully populated RETE net $(N^{\Phi}_{Q},p^{\Phi}_{Q})$, it must hold that $\sum_{n^{\Phi}_{Q^{\prime}}\in V^{N^{\Phi}_{Q^{\prime}}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime}})|\leq 10\cdot\sum_{n_{Q^{\prime}}\in V^{N_{Q^{\prime}}}}|\mathcal{C}(n_{Q^{\prime}})|$. Also, by the semantics of the marking-sensitive semi-join and because the result set for $p^{\Phi}_{Q^{\prime}}$ contains each match for $Q^{\prime}$ at most once by Lemma 12 in [BG24b], it must hold that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi})}|m|\leq\sum_{m\in\mathcal{C}(p_{Q^{\prime}})}|m|\leq\sum_{n_{Q^{\prime}}\in V^{N_{Q^{\prime}}}}|\mathcal{C}(n_{Q^{\prime}})|$.
Any nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$ must have some nested query $(Q^{\prime\prime},\exists(a:Q^{\prime\prime}\rightarrow Q^{\prime\prime\prime},\psi^{\prime\prime}))$ or $(Q^{\prime\prime},\texttt{true})$ in order for $\psi$ by the definition of nested graph conditions. In either case, $N$ has to contain a corresponding subnet $(N_{Q^{\prime\prime}},p_{Q^{\prime\prime}})$ computing matches for $Q^{\prime\prime}$. It then holds for this subnet $N_{Q^{\prime\prime}}$ by Theorem 5 that $\sum_{n^{\Phi}_{Q^{\prime\prime}}\in V^{N^{\Phi}_{Q^{\prime\prime}}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime\prime}})|\leq 10\cdot\sum_{n_{Q^{\prime\prime}}\in V^{N_{Q^{\prime\prime}}}}|\mathcal{C}(n_{Q^{\prime\prime}})|$. It also must hold that $|\mathcal{C}^{\Phi}([\cup]^{\Phi})|\leq|\mathcal{C}(p_{Q^{\prime\prime}})|\leq\sum_{n_{Q^{\prime\prime}}\in V^{N_{Q^{\prime\prime}}}}|\mathcal{C}(n_{Q^{\prime\prime}})|$ and $\sum_{n^{\Phi}\in V^{RPS_{l}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 3\cdot\sum_{n_{Q^{\prime\prime}}\in V^{N_{Q^{\prime\prime}}}}|\mathcal{C}(n_{Q^{\prime\prime}})|$.
@@
In the worst case, it must thus hold that $\sum_{n^{sat}\in V^{N^{sat}}}\mathcal{C}^{\Phi}(n^{sat})|\leq 25\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $(X^{\Delta},p_{X}^{\Delta})$ a modular extension of the marking-sensitive RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for $(X^{\Delta},p_{X}^{\Delta})$. Executing $(X^{\Delta},p_{X}^{\Delta})$ via $O=order^{\Delta}(N^{\Delta})$ then yields a configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Delta},H,H_{p},\mathcal{C}^{\Phi}_{0})$ that is consistent for all nodes in $N^{\Delta}$.
@@
Follows directly from Lemma 13 in [BG24b] and Lemmata 45 and 40.

### v2

@@ A.5. Supplementary Lemmata
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ consists of $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$, $(N^{sat}_{(Q^{\prime},\psi^{\prime})},p^{sat}_{(Q^{\prime},\psi^{\prime})})=localize^{sat}(Q^{\prime},\psi^{\prime})$, $[\cup]^{\Phi}$, $RPS^{\infty}_{r}=RPS([\cup]^{\Phi},N^{\Phi}_{Q})$, and $p^{sat}=[\ltimes]^{\Phi}$. The execution order for $(N^{sat},p^{sat})$ is then given by $order^{sat}(N^{sat})=order(N^{\Phi}_{Q^{\prime}})\circ order^{sat}(N^{sat}_{(Q^{\prime},\psi^{\prime})})\circ[\cup]^{\Phi}\circ toposort(RPS^{\infty}_{r})^{-1}\circ order(N^{\Phi}_{Q})\circ[\ltimes]^{\Phi}$.
By Lemma 13 in [BG24b], $\mathcal{C}_{1}^{\Phi}=execute(order(N^{\Phi}_{Q^{\prime}}),N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$ is consistent for all nodes in $N^{\Phi}_{Q^{\prime}}$. $\mathcal{C}_{2}^{\Phi}=execute(order^{sat}(N^{sat}_{(Q^{\prime},\psi^{\prime})}),N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{1})$ is then also consistent for all nodes in $N^{sat}_{Q^{\prime}}$ by the induction hypothesis. By Lemma 13 in [BG24b], $\mathcal{C}^{\Phi}=execute([\cup]^{\Phi}\circ toposort(RPS^{\infty}_{r})^{-1}\circ order(N^{\Phi}_{Q})\circ[\ltimes]^{\Phi},N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{2})$ must then be consistent for all nodes in $N^{sat}$.
@@
(RETE net localization via $localize^{sat}$ introduces an overhead of at most factor 25 on the number of computed matches.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. It then holds that $\sum_{n^{sat}\in V^{N^{sat}}}|\mathcal{C}^{\Phi}(n^{sat})|\leq 25\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@
For each nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$, $N^{sat}$ contains the subnets $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$ and $(N^{\Phi}_{Q^{\prime\prime}},p^{\Phi}_{Q^{\prime\prime}})=localize(Q^{\prime\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$, a marking-sensitive semi-join $[\ltimes]^{\Phi}$, as well as the request projection structure $RPS^{\infty}_{l}=RPS^{\infty}([\cup]^{\Phi},N^{\Phi}_{Q^{\prime}})$. $N$ contains a RETE subnet $(N_{Q^{\prime}},p_{Q^{\prime}})$ computing matches for $Q^{\prime}$ and a semi-join []delimited-[]left-normal-factor-semidirect-product[$\ltimes$][ ⋉ ].
By Lemma 26 and because the additional dependency on $RPS^{\infty}_{l}$ does not affect the worst case of a fully populated RETE net $(N^{\Phi}_{Q},p^{\Phi}_{Q})$, it must hold that $\sum_{n^{\Phi}_{Q^{\prime}}\in V^{N^{\Phi}_{Q^{\prime}}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime}})|\leq 10\cdot\sum_{n_{Q^{\prime}}\in V^{N_{Q^{\prime}}}}|\mathcal{C}(n_{Q^{\prime}})|$. Also, by the semantics of the marking-sensitive semi-join and because the result set for $p^{\Phi}_{Q^{\prime}}$ contains each match for $Q^{\prime}$ at most once by Lemma 12 in [BG24b], it must hold that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi})}|m|\leq\sum_{m\in\mathcal{C}(p_{Q^{\prime}})}|m|\leq\sum_{n_{Q^{\prime}}\in V^{N_{Q^{\prime}}}}|\mathcal{C}(n_{Q^{\prime}})|$.
Any nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$ must have some nested query $(Q^{\prime\prime},\exists(a:Q^{\prime\prime}\rightarrow Q^{\prime\prime\prime},\psi^{\prime\prime}))$ or $(Q^{\prime\prime},\texttt{true})$ in order for $\psi$ by the definition of nested graph conditions. In either case, $N$ has to contain a corresponding subnet $(N_{Q^{\prime\prime}},p_{Q^{\prime\prime}})$ computing matches for $Q^{\prime\prime}$. It then holds for this subnet $N_{Q^{\prime\prime}}$ by Theorem 5 that $\sum_{n^{\Phi}_{Q^{\prime\prime}}\in V^{N^{\Phi}_{Q^{\prime\prime}}}}|\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime\prime}})|\leq 10\cdot\sum_{n_{Q^{\prime\prime}}\in V^{N_{Q^{\prime\prime}}}}|\mathcal{C}(n_{Q^{\prime\prime}})|$. It also must hold that $|\mathcal{C}^{\Phi}([\cup]^{\Phi})|\leq|\mathcal{C}(p_{Q^{\prime\prime}})|\leq\sum_{n_{Q^{\prime\prime}}\in V^{N_{Q^{\prime\prime}}}}|\mathcal{C}(n_{Q^{\prime\prime}})|$ and $\sum_{n^{\Phi}\in V^{RPS^{\infty}_{l}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 3\cdot\sum_{n_{Q^{\prime\prime}}\in V^{N_{Q^{\prime\prime}}}}|\mathcal{C}(n_{Q^{\prime\prime}})|$.
@@
In the worst case, it must thus hold that $\sum_{n^{sat}\in V^{N^{sat}}}|\mathcal{C}^{\Phi}(n^{sat})|\leq 25\cdot\sum_{n\in V^{N}}|\mathcal{C}(n)|$.
@@
(Execution of modular extensions of localized RETE nets via $order^{\Delta}$ yields consistent configurations for the base RETE net.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $(X^{\Delta},p_{X}^{\Delta})$ a modular extension of the marking-sensitive RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for $(X^{\Delta},p_{X}^{\Delta})$. Executing $(X^{\Delta},p_{X}^{\Delta})$ via $O=order^{\Delta}(N^{\Delta})$ then yields a configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Delta},H,H_{p},\mathcal{C}^{\Phi}_{0})$ that is consistent for all nodes in $N^{\Delta}$.
@@
Follows directly from Lemma 13 in [BG24b] and Lemmata 33 and 28.
Appendix B Queries
Figure 24. Simple path query for the synthetic abstract syntax graph scenario
Figure 25. Composite query for the real abstract syntax graph scenario
Figure 26. Observer query for the real abstract syntax graph scenario
Figure 27. Singleton query for the real abstract syntax graph scenario
Figure 28. Strategy query for the real abstract syntax graph scenario
Figure 29. Query “Interactive 1” for the LDBC scenario
Figure 30. Query “Interactive 2” for the LDBC scenario
Figure 31. Query “Interactive 3” for the LDBC scenario
Figure 32. Query “Interactive 4” for the LDBC scenario
Figure 33. Query “Interactive 5” for the LDBC scenario
Figure 34. Query “Interactive 6” for the LDBC scenario
Figure 35. Query “Interactive 7” for the LDBC scenario
Figure 36. Query “Interactive 8” for the LDBC scenario
Figure 37. Query “Interactive 9” for the LDBC scenario
Figure 38. Query “Interactive 10” for the LDBC scenario
Figure 39. Query “Interactive 11” for the LDBC scenario
Figure 40. Query “Interactive 12” for the LDBC scenario
Figure 41. Query “Interactive 3 NGC” for the LDBC scenario. The query includes a nested graph condition of the form $\neg\exists(a_{1}:Q\rightarrow Q_{1}^{\prime},\texttt{true})\wedge\neg\exists(a_{2}:Q\rightarrow Q_{2}^{\prime},\texttt{true})$, with the parts of $Q_{1}^{\prime}$ and $Q_{2}^{\prime}$ that are not in the image of $a_{1}$ respectively $a_{2}$ inside the boxes labeled “NAC1” and “NAC2”
Figure 42. Query “Interactive 4 NGC” for the LDBC scenario. The query includes a nested graph condition of the form $\neg\exists(a:Q\rightarrow Q^{\prime},\texttt{true})$, with the part of $Q^{\prime}$ that is not in the image of $a$ inside the box labeled “NAC”